OpenStudy (anonymous):

Solve the system using elimination. 3x + 3y = –9 3x – 3y = 21

OpenStudy (anonymous):

add the two equation, the y variable will be eliminated

OpenStudy (anonymous):

then solve for x

OpenStudy (anonymous):

once you get what x is, you can replace it into one of the original equation to get y

OpenStudy (anonymous):

can you give me an example

OpenStudy (anonymous):

3x + 3y = –9 (1) 3x – 3y = 21 (2) So: (1) + (2) => 6x=12=>x=2=>y=-5. So (2,-5) is right